### v1

# A preconditioned difference of convex functions algorithm with extrapolation and line search

Abstract
This paper proposes a novel proximal difference-of-convex (DC) algorithm enhanced with extrapolation and aggressive non-monotone line search for solving non-convex optimization problems. We introduce an adaptive conservative update strategy of the extrapolation parameter determined by a computationally efficient non-monotone line search. The core of our algorithm is to unite the update of the extrapolation parameter with the step size of the non-monotone line search interactively. The global convergence of the two proposed algorithms is established through the Kurdyka-Łojasiewicz properties, ensuring convergence within a preconditioned framework for linear equations. Numerical experiments on two general non-convex problems: SCAD-penalized binary classification and graph-based Ginzburg-Landau image segmentation models, demonstrate the proposed method’s high efficiency compared to existing DC algorithms both in convergence rate and solution accuracy.
Key words.
difference of convex functions, extrapolation, preconditioning, nonmonotone line search, Kurdyka-Łojasiewicz properties, global convergence.
MSCodes.
65K10, 65F08, 49K35, 90C25, 90C26.
1 Introduction
We aim to solve the following non-convex optimization problem
$\min_{x\in X}\ E(x):=f(x)+g(x)\,=\,f\left(x\right)+g_{1}\left(x\right)-g_{2}\left(x\right).$ (1.1)
Here $g_{1}\left(x\right)$ and $g_{2}\left(x\right)$ are proper closed convex functions, $\nabla f\left(x\right)$ is Lipschitz continuous with constants $L$, and $X$ is a finite-dimensional Hilbert space. Such non-convex optimization formulations find broad applications in digital image processing, statistical machine learning, and related fields [1, 8, 11]. The proximal DC algorithm (DCA) with extrapolation (shortened as pDCAe henceforth) has emerged as a particularly efficient approach, successfully accelerating conventional proximal gradient methods [18, 17, 19]. This method combines the computational efficiency of proximal gradient algorithms with DC decomposition techniques, as detailed in [1, 11] for various non-convex optimizations. Our work builds upon these foundations while introducing critical enhancements to the algorithmic framework.
As shown in the survey [19, Section 5], the convergence speed of DCA can be accelerated through two distinct approaches: extrapolation techniques with traditional parameterization strategies, including FISTA-based extrapolation [25, Algorithm 1] or FISTA-based restart mechanisms, e.g., [19, Algorithm 5], [29, Algorithm (3.4)] and Armijo line search methods based on energy decreasing [2, Algorithm 2], [27].
Compared to the existing works, our contributions are as follows. The main contributions are a concise, lightweight non-monotone line search mechanism and a new adaptive update strategy for the extrapolation parameter. The line search completely determines the extrapolation parameter. Inspired by [2, 3, 15], we employ the aggressive non-monotone line search, where the starting point for line search is generated by the preconditioned and extrapolated DCA, instead of using the previous iteration [2, 3]. The aggressive non-monotone line search [15, Example 4] can not only relax stringent descent requirements, but can also obtain an extra step. The proposed framework enhances pDCAe by integrating these two accelerating methods. Specifically, the non-monotone line search strategy provides global guidance for local extrapolation parameter update. We refer to [31, NLSA] [23, Algorithm 1] [12, Section 4] for the traditional non-monotone line search strategy. Moreover, given the limited existing work that combines these approaches, our comparative analysis demonstrates distinct advantages over existing methods. Compared to adaptive linear search methods [30, Algorithm 1] that require the solving of auxiliary subproblems, our approach eliminates computational overhead while preserving the core efficiency of pDCAe.
Furthermore, we introduce a line search framework for the case $g_{1}$ being non-smooth under new and mild conditions with a global convergence guarantee, which is not covered by the line search framework [3]. In [3], only the case $g_{2}$ is non-smooth, and $g_{1}$ with Lipschitz gradient is discussed to ensure the existence of descent directions.
Numerical validations confirm the improvements. In Section 5 for numerical experiments, we compare our algorithm with the algorithms that use only a single type of acceleration method. The experiments show that our algorithm has better performance. First, for iterative sequences near critical points, the integrated update strategy stabilizes convergence by controlling the extrapolation parameter $\beta$, effectively addressing the performance degradation observed in the original implementations of pDCAe. Second, the interactive combination of extrapolation and line search accelerates convergence compared to standard extrapolation and restart approaches. These empirical results, detailed in Section 5, verify that our framework maintains the computational simplicity of pDCAe while achieving better convergence across various non-convex optimization benchmarks.
The paper is organized as follows. Section 2 establishes notations and preliminary materials essential for the subsequent analysis. Section 3 introduces two distinct DC decomposition algorithms incorporating a non-monotone line search strategy and a novel update strategy of the extrapolation parameter. Section 4 rigorously analyzes the convergence properties of the proposed algorithms and examines a generalized computational framework. Finally, Section 5 validates the theoretical developments through numerical experiments, and Section 6 gives a conclusion.
2 Notation and preliminaries
[…]
We will touch on some necessary tools from the convex and variational analysis [26]. The graph of a multi-valued mapping $F:\mathbb{R}^{n}{\rightrightarrows}\mathbb{R}^{m}$ is defined by
$\operatorname{gph}F:=\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{m}:y\in F(x)\},$
whose domain is defined by $\operatorname{dom}F:=\{x\ |\ F(x)\neq\emptyset\}$. Similarly the graph of an extended real-valued function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined by
$\operatorname{gph}f:=\{(x,s)\in\mathbb{R}^{n}\times\mathbb{R}:s=f(x)\}.$
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Denote $\operatorname{dom}h:=\{x\in\mathbb{R}^{n}:\ h(x)<+\infty\}$. For each $x\in\operatorname{dom}f$, the limiting-subdifferential of $h$ at $x\in\mathbb{R}^{n}$, written $\partial f$, is defined as follows [26],
[…]
We next recall the Kurdyka-Łojasiewicz (KL) property, which is satisfied by a wide variety of functions, such as proper closed semi-algebraic functions. It plays an important role in the convergence analysis of many first-order methods.
Let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper lower semicontinuous function. Denote $\operatorname{dom}h:=\{x\in\mathbb{R}^{n}:\ h(x)<+\infty\}$. For each $x\in\operatorname{dom}f$, the limiting-subdifferential of $h$ at $x\in\mathbb{R}^{n}$, written $\partial f$ are well-defined as in [26, Definition 8.3(a)]. Additionally, if $h$ is continuously differentiable, the subdifferential reduces to the usual gradient $\nabla h$. We also need the following Kurdyka-Łojasiewicz (KL) property and KL exponent for the global and local convergence analysis. While the KL properties can help obtain the global convergence of iterative sequences, the KL exponent can help provide a local convergence rate.
Definition 1 (KL property, KL function and KL exponent [4]).
A proper closed function $h$ is said to satisfy the KL property at $\bar{x}\in\operatorname{dom}\partial h$ if there exists $\nu\in(0,+\infty]$, a neighborhood $\mathcal{O}$ of $\bar{x}$, and a continuous concave function $\psi:[0,\nu)\rightarrow[0,+\infty)$ with $\psi(0)=0$ such that
  • (i)
    $\psi$ is continuous differentiable on $(0,\nu)$ with $\psi^{\prime}>0$ over $(0,\nu)$;
  • (ii)
    for any $x\in\mathcal{O}$ with $h(\bar{x})<h(x)<h(\bar{x})+\nu$, one has
    $\psi^{\prime}(h(x)-h(\bar{x}))\operatorname{dist}(0,\partial h(x))\geq 1.$ (2.1)
A proper closed function $h$ satisfying the KL property at all points in $\operatorname{dom}\partial h$ is called a KL function. Furthermore, for a proper closed function $h$ satisfying the KL property at $\bar{x}\in\operatorname{dom}\partial h$, if $\psi$ in (2.1) can be chosen as $\psi(s)=cs^{1-\theta}$ for some $\theta\in[0,1)$ and $c>0$, i.e., there exist $\bar{c},\epsilon>0$ such that
$\operatorname{dist}(0,\partial h(x))\geq\bar{c}(h(x)-h(\bar{x}))^{\theta}$ (2.2)
whenever $\|x-\bar{x}\|\leq\epsilon$ and $h(\bar{x})<h(x)<h(\bar{x})+\nu$, then we say that $h$ has the KL property at $\bar{x}$ with exponent $\theta$. If $h$ has the KL property with exponent $\theta$ at any $\bar{x}\in\operatorname{dom}\partial h$, we call $h$ a KL function with exponent $\theta$.
[…]
3 Preconditioned DCA with extrapolation determined by line search
Now let us turn to preconditioned DCA with extrapolation for (1.1). By solving the subproblem involving the difference of convex functions by incorporating a proximal term to update $\bar{x}^{n}$, the optimization task for determining $\bar{x}^{n}$ is as follows
$\bar{x}^{n}=\mathop{\operatorname*{arg\,min}}\limits_{y\in\mathbb{R}^{n}}\left\{\left<-\xi^{n},y\right>+\frac{1}{2}\left\|y-y^{n}\right\|_{M}^{2}+f\left(y\right)+g_{1}\left(y\right)\right\},$ (3.1)
where $\xi^{n}\in\partial g_{2}\left(x^{n}\right)$, $\forall n\in N$, and $y^{n}=x^{n}+\beta_{n}\left(x^{n}-x^{n-1}\right)$.
Henceforth, we assume that the symmetric positive definite matrix $M$ serves as a preconditioning weight operator, where $\langle a,b\rangle_{M}=\langle a,Mb\rangle_{M}$ for $\forall$ $a,b\in X$ and the induced norm as $\left\|x\right\|_{M}^{2}:=\left<Mx,x\right>$. Applying first-order optimality conditions to the minimization problem (3.1) yields the critical equation for determining $\bar{x}^{n}$
$-\xi^{n}+M\left(y-y^{n}\right)+\nabla f\left(y\right)+\partial g_{1}\left(y\right)=0.$ (3.2)
We assume the above equation admits a unique solution, yielding the following update
$\bar{x}^{n}=\left(M+\nabla f+\partial g_{1}\right)^{-1}\left(My^{n}+\xi^{n}\right).$ (3.3)
If $g_{1}+f$ is a quadratic function, the update $\bar{x}^{n}$ in (3.3) can be reformulated as the classical efficient preconditioned iteration, which can circumvent solving the linear equation even with mild accuracy, especially for a large-scale linear system. We will explain the preconditioning technique in Proposition 2 later.
Within the widely employed gradient descent framework, [24] pioneered two distinct restart strategies for extrapolation parameter: a gradient-based scheme [16, 14, 29] and a function-value-based scheme. The former, taking [29] for example, represents a computationally efficient implementation, while the latter, a function value-based restart scheme, i.e., restarting whence $E\left(x^{n}\right)>E\left(x^{n-1}\right)$, which conceptually parallels the line search techniques proposed here, provides enhanced theoretical guarantees. We also refer to [6] for other restart techniques with the estimation of Lipschitz constants based on the FISTA framework. However, the function-value-based restart scheme usually needs to be restarted frequently due to the strict energy-decreasing requirement, which leads to an increased computational cost, as observed in our numerical experiments. Crucially, we address a critical failure mode: persistent line search failures triggering an excessive restart, which degrade the algorithm to basic DCA performance. To mitigate this, we directly introduce a novel update strategy that dynamically adjusts the extrapolation parameter based on the following line search. By defining $d^{n}=\bar{x}^{n}-x^{n}$, we proceed with an aggressive non-monotone Armijo-type line search involving parameters $\rho^{k}\lambda_{\max}>0$, $k=1,2,\ldots$, $\rho\in(0,1)$, and $\lambda_{\max},\omega,\eta>0$. The aim is to find the smallest integer $k$ satisfying [15]
$E(\bar{x}^{n}+\rho^{k-1}\lambda_{\max}d^{n})\leq{E}(\bar{x}^{n})-\eta\rho^{k-1}\lambda_{\max}\|d^{n}\|_{2}^{2}+\nu_{n},\quad\nu_{n}:=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}.$ (3.4)
Henceforth, denote $a(n):=k$ as the number of line searches for the $n$-th iteration and $N_{\max}$ as the maximal number of line searches. The final step size $\lambda_{n}:=\rho^{k-1}\lambda_{\max}$ obtained from the line search in (3.4) dictates the update of $x^{n}$ as follows
$x^{n+1}=\bar{x}^{n}+{\lambda}_{n}d^{n}.$ (3.5)
We propose the following method to update the extrapolation parameter $\beta_{n+1}$ determined by line search (3.4). We assume that $b_{1}>0$ and $b_{2}\geq 0$ are constants close to 0 henceforth.
Line search determined extrapolation (shortened as LSDE henceforth):
[…]
A critical distinction between our algorithm and conventional FISTA-based methods lies in parameterizing the extrapolation update $\left\{\beta_{n+1}\right\}$. Established schemes-including FISTA with fixed or adaptive restart-initialize $\theta_{-1}=\theta_{0}=1$ and recursively compute $\beta_{n+1}$ as
$\beta_{n+1}=\frac{\theta_{n}-1}{\theta_{n+1}},\quad\theta_{n+1}=\frac{1+\sqrt{1+4\theta_{n}^{2}}}{2},$ (3.7)
with parameter resets $\theta_{n}=\theta_{n+1}=1$ under specific conditions: fixed restart triggers resets periodically every $\bar{T}$ iterations, whereas adaptive restart activates when $y^{n}-x^{n+1},x^{n+1}-x^{n}\right>>0$.
Proposition 1.
According to (3.6) for updating the extrapolation parameter $\beta_{n+1}$, we assert that there exists a positive constant $C_{\lambda}$ such that $\frac{1}{\left(1+\lambda_{n}\right)^{2}}-\beta_{n+1}^{2}>C_{\lambda}$ for $\forall n$.
Proof.
Recalling that $\frac{1}{\left(1+\lambda_{n}\right)^{2}}>\beta_{n+1}^{2}$, we then discuss this in two scenarios.
If $\lambda_{n}\neq 0$, known that $\beta_{n+1}=\frac{1}{1+b_{1}+\lambda_{n}}$, we then obtain
$\frac{1}{\left(1+\lambda_{n}\right)^{2}}-\frac{1}{\left(1+b_{1}+\lambda_{n}\right)^{2}}=\frac{b_{1}\left(1+b_{1}+2\lambda_{n}\right)}{\left(1+\lambda_{n}\right)^{2}\left(1+b_{1}+\lambda_{n}\right)^{2}}\geq\frac{b_{1}\left(1+b_{1}\right)}{\left(1+\lambda_{\max}\right)^{2}\left(1+b_{1}+\lambda_{\max}\right)^{2}}\triangleq C_{\lambda}^{1}.$
If $\lambda_{n}=0$, we know that $\beta_{n+1}=0$, thus we have
$\frac{1}{\left(1+\lambda_{n}\right)^{2}}-\beta_{n+1}^{2}\geq 1-\varepsilon_{0}>C_{\lambda}^{2}.$
Setting $C_{\lambda}=\min\left(C_{\lambda}^{1},C_{\lambda}^{2}\right)$, we proved this conclusion. ∎
Specifically, we introduce two distinct DC splitting algorithms, both integrated with a non-monotone line search strategy and our proposed update strategy (3.6) for solving the minimization problem (1.1); see Algorithm 1 and Algorithm 2 for more details. If putting $f(y)$ in the implicit part as $g_{1}$, we get Algorithm 1. While putting $f(y)$ in the explicit form with linearization as $g_{2}$ [29], we obtain Algorithm 2.
[…]
1:Choose $x^{0}\in\operatorname{dom}g_{1},\left\{\beta_{n}\right\}\subset\left[0,1\right)$ and $\mathop{\sup}\limits_{n}\beta_{n}<1$, $x^{-1}=x^{0}$, $\beta_{0}=0$.
2:Choose $\xi^{n}\in\partial g_{2}\left(x^{n}\right),\forall n\in N$, $y^{n}=x^{n}+\beta_{n}\left(x^{n}-x^{n-1}\right)$.
3:Solve the following subproblem
$\bar{x}^{n}=\mathop{\operatorname*{arg\,min}}\limits_{y\in X}\left\{\left<-\xi^{n},y\right>+\frac{1}{2}\left\|y-y^{n}\right\|^{2}_{M}+f\left(y\right)+g_{1}\left(y\right)\right\}.$ (3.8)
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$. Otherwise, go to Step 8.
5:Set $\lambda_{\max}>0$, $\eta>0$, $\rho\in\left[0,1\right)$, $N_{\max}\in\mathbb{N}$.
6:Set $\nu_{n}=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}$, $a\left(n\right)=1$, $\lambda_{n}=\lambda_{\max}$.
7:Start the non-monotone line search
IF $\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\ \lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}.$ (3.9)
8:Do $\lambda_{n}=\rho\cdot\lambda_{n}$, set $a\left(n\right)=a\left(n\right)+1$; Otherwise, RETURN $x^{n+1}:=\bar{x}^{n}+\lambda_{n}d^{n}$.
9:If $a\left(n\right)>N_{\max}$, STOP and RETURN $x^{n+1}:=\bar{x}^{n}$; Otherwise, go to Step 8.
10:Update $\beta_{n+1}$ as LSDE (3.6).
[…]
1:Choose $x^{0}\in\operatorname{dom}g_{1},\left\{\beta_{n}\right\}\subset\left[0,1\right)$ and $\mathop{\sup}\limits_{n}\beta_{n}<1$, $x^{-1}=x^{0}$, $\beta_{0}=0$.
2:Choose $\xi^{n}\in\partial g_{2}\left(x^{n}\right),\forall n\in N$, $y^{n}=x^{n}+\beta_{n}\left(x^{n}-x^{n-1}\right)$.
3:Solve the following subproblem
$\bar{x}^{n}=\mathop{\operatorname*{arg\,min}}\limits_{y\in X}\left\{\left<\nabla f\left(y^{n}\right)-\xi^{n},y\right>+\frac{L}{2}\left\|y-y^{n}\right\|_{M}^{2}+g_{1}\left(y\right)\right\}.$ (3.10)
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$. Otherwise, go to Step 8.
5:Set $\lambda_{\max}>0$, $\eta>0$, $\rho\in\left[0,1\right)$, $N_{\max}\in\mathbb{N}$.
6:Set $\nu_{n}=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}$, $a\left(n\right)=1$, $\lambda_{n}=\lambda_{\max}$.
7:Start the non-monotone line search
IF $\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\eta\lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}.$ (3.11)
8:Do $\lambda_{n}=\rho\cdot\lambda_{n}$, set $a\left(n\right)=a\left(n\right)+1$; Otherwise, RETURN $x^{n+1}:=\bar{x}^{n}+\lambda_{n}d^{n}$.
9:If $a\left(n\right)>N_{\max}$, STOP and RETURN $x^{n+1}:=\bar{x}^{n}$; Otherwise, go to Step 8.
10:Update $\beta_{n+1}$ as LSDE (3.6).
To better show the novelty of our algorithms, we collect some commonly used DC algorithms and their respective characteristics in the following Table 1.
Table 1: Commonly used DC algorithms and their respective frameworks.
accelerated approach extrapolation line search preconditioner algorithm
I FISTA (restart) No No pDCAe [29]
II No monotone No BDCA [2]
III No non-monotone No nmBDCA [15]
IV FISTA (restart) No Yes preDCAe [13]
V LSDE non-monotone Yes Our work
Before starting the proof of convergence, we shall give the strategy [15] to choose the parameter sequence $\nu_{n}$.
Remark 1.
Let $\omega>0$ be a constant, and the sequence $\left\{\nu_{n}\right\}_{n\in N}$ defined by $\nu_{n}=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}$, for all $n\in N$. Indeed, due to $\lim_{n\rightarrow\infty}\nu_{n}=\lim_{n\rightarrow\infty}\frac{\omega}{n+1}=0$, for every $\delta>0$, there exists $n_{0}\in N$ such that $n\geq n_{0}$ implies $\frac{\omega}{n+1}\leq\delta<\eta\cdot\lambda_{\max}\cdot\rho^{N_{\max}}$. We then have $\nu_{n}\leq\delta\left\|d^{n}\right\|^{2}$. Similarly, we can define $\nu_{n}=\frac{\omega}{p\left(n\right)}\left\|d^{n}\right\|^{2}$, where $p\left(n\right)$ is a monotonically increasing function concerning $n$, here the choice of $p\left(n\right)$ depends on the rate at which we want the non-monotone term to decline.
This brings great flexibility. At the beginning of iterations, the line search (3.4) with $\nu_{n}$ is indeed non-monotone. However, it eventually becomes monotone for sufficiently large $n$ since $\nu_{n}$ converges to 00 while $N_{\max}$ is bounded.
Now, we end this section with the following proposition for addressing the preconditioning.
Proposition 2.
If $g_{1}+f$ is quadratic function with $\nabla(g_{1}+f)(y)=Ay-b_{0}$ and $A$ is a positive definite and linear operator $A$, the update in (3.3) can be reformulated as the following preconditioned iteration
$\bar{x}^{n}=y^{n}+\mathbb{M}^{-1}(b^{n}-Ay^{n}),$ (3.12)
where
$\mathbb{M}=A+M,\quad b^{n}=b_{0}+\xi^{n},$
which is one time classical preconditioned iteration for solving (3.2), i.e., $Ay=b^{n}$ with initial value $y^{n}$.
Proof.
With the assumptions, the equation (3.2) becomes
$-\xi^{n}+(M+A)y-My^{n}-b_{0}=0,$
which is equivalent to
$y$ $=(M+A)^{-1}(My^{n}+\xi^{n}+b_{0})=(M+A)^{-1}((M+A)y^{n}+\xi^{n}+b_{0}-Ay^{n})$
$=y^{n}+(M+A)^{-1}(\xi^{n}+b_{0}-Ay^{n})=y^{n}+(M+A)^{-1}(b^{n}-Ay^{n}).$
By letting $\bar{x}^{n}=y$, the proof is finished. We emphasize that the preconditioned iteration automatically determines $M$, and no explicit expression is needed. We refer to [9, Section 2.3] for more details and [28, Proposition 3.2] for a similar idea. ∎
4 Convergence analysis
This section analyzes the convergence of the two proposed algorithms and addresses a generalized special case. We first present the convergence for the case that $f(x)$ and $g_{1}(x)$ have Lipschitz gradients while $g_{2}$ is non-smooth. For the case $g_{1}$ being non-smooth, we rigorously prove convergence for the scenario where $g_{1}(x)$ is non-differentiable and $g_{2}(x)$ is Lipschitz differentiable under mild condition. This proof imposes standard assumptions on and selects a representative special case to demonstrate the core methodology. Collectively, these two cases can cover the most practical computational scenarios encountered in applications.
4.1 Convergence analysis of Algorithm 1
We now begin our convergence analysis by establishing the following lemma that illustrates the quasi-descent property inherent in the underlying energy function.
Lemma 1.
Let $\bar{x}^{n}$ is generated by (3.8), then the following inequality holds
$E\left(\bar{x}^{n}\right)\leq E\left(x^{n}\right)+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$ (4.1)
Proof.
As $\bar{x}^{n}$ is a minimizer of the strongly convex function in (3.8) and by evaluating the function values at $\bar{x}^{n}$ and $x^{n}$, we obtain
$-\xi^{n},\bar{x}^{n}\right>+\frac{1}{2}\left\|\bar{x}-y^{n}\right\|_{M}^{2}+f\left(\bar{x}^{n}\right)+g_{1}\left(\bar{x}^{n}\right)\leq\left<-\xi^{n},x^{n}\right>+\frac{1}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
$+f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$ (4.2)
Given $\xi^{n}\in\partial g_{2}\left(x^{n}\right)$, the convexity of $g_{2}$ implies
$g_{2}\left(\bar{x}^{n}\right)\geq g_{2}\left(x^{n}\right)+\left<\xi^{n},\bar{x}^{n}-x^{n}\right>\Rightarrow\langle-\xi^{n},\bar{x}^{n}-x^{n}\rangle\geq g_{2}(x^{n})-g_{2}(\bar{x}^{n}).$
Substituting $\langle-\xi^{n},\bar{x}^{n}-x^{n}\rangle$ with $g_{2}(x^{n})-g_{2}(\bar{x}^{n})$ into the left-hand side of (4.2) yields
$g_{2}\left(x^{n}\right)-g_{2}\left(\bar{x}^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+f\left(\bar{x}^{n}\right)+g_{1}\left(\bar{x}^{n}\right)\leq\frac{1}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
$+f\left(x^{n}\right)+g\left(x^{n}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Rearranging the above terms, we obtain
$f\left(\bar{x}^{n}\right)+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(\bar{x}^{n}\right)$ $\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)+\frac{1}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
$-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Remembering $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, this leads to the inequality (4.1). ∎
This important descent characteristic rigorously justifies the adoption of a non-monotone line search strategy. Below, we provide a concise characterization of this mechanism’s mathematical foundation, verifying its well-posedness through the following systematic analysis.
Remark 2.
The relationship between non-monotone and monotone line search in Algorithm 1 is governed by the parameter $\nu_{n}$. When $\nu_{n}=0$ for all iterations $n$, the non-monotone line search reduces exactly to its monotone counterpart, requiring strict energy descent. However, the non-monotone line search also requires the critical descent condition
$\nu_{n}<\eta\cdot\lambda_{n}\cdot\left\|d^{n}\right\|^{2},\ \text{for sufficiently large}\ n.$
This inequality is readily satisfiable under our parameterization of $\nu_{n}$. Specifically, for $n>n_{0}$ (where $n_{0}$ is problem-dependent), we obtain
$\frac{w}{n+1}<\eta\cdot\lambda_{\max}\cdot\rho^{N_{\max}}.$
Here $\lambda_{\text{max}}$ denotes the maximum admissible step size and $\rho^{N_{\text{max}}}$ reflects the worst-case step size contraction. While the line search permits temporary function value increases during initial iterations (enabled by $\nu_{n}$), the descent dominance condition $\nu_{n}<\eta\cdot\lambda_{n}\cdot\left\|d^{n}\right\|^{2}$, guarantees eventual monotone convergence. Furthermore, this relaxation enables temporary increases in the function value while maintaining global convergence, which is a critical feature for escaping shallow local minima in non-convex landscapes [3, Example 3.4].
We now establish global convergence properties for the proposed Algorithm 1. This result will serve as the foundation for deriving global convergence of the sequence $x^{n}$ under appropriate regularity conditions.
Lemma 2.
Let $\left\{x^{n}\right\}_{n}$ is generated from $npDCAe_{nls}$ as Algorithm 1 for solving (1.1), then we have
$\sum_{n=1}^{\infty}{\left\|x^{n+1}-x^{n}\right\|^{2}<\infty}.$ (4.3)
Proof.
Building upon the inequality (4.1), we now incorporate it with the non-monotone line search. When a suitable step size $\lambda_{n}>0$ in Case 1 of (3.6) satisfies the search criterion, we obtain
$E\left(x^{n+1}\right)\triangleq E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)\leq E\left(\bar{x}^{n}\right)-\eta\cdot\lambda_{n}\cdot\left\|d^{n}\right\|^{2}+\nu_{n}.$
Combining these results yields
$E\left(x^{n+1}\right)\leq E\left(x^{n}\right)-\eta\cdot\lambda_{n}\cdot\left\|d^{n}\right\|^{2}-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}+\nu_{n}.$
Rearranging terms produces the key inequality
$\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}$
$\leq E\left(x^{n}\right)-E\left(x^{n+1}\right)=f\left(x^{n}\right)+g\left(x^{n}\right)-f\left(x^{n+1}\right)-g\left(x^{n+1}\right).$ (4.4)
For Case 2 of (3.6), while the line search does not succeed with $N_{\max}$ searches, we assume $\lambda_{n}=0$ along with $\nu_{n}$ set to 0 and hence $x^{n+1}=\bar{x}^{n}$ as in (4.1), i.e.,
$E\left({x}^{n+1}\right)\leq E\left(x^{n}\right)+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{1}{2}\left\|{x}^{n+1}-x^{n}\right\|_{M}^{2}.$ (4.5)
For unifying the analysis of both cases, we first define an indicator function
$\delta_{\lambda}^{n}=\begin{cases}1\quad\lambda_{n}>0\\ 0\quad\lambda_{n}=0\\ \end{cases}.$
We now combine the inequality (4.4) and (4.5) with $\delta_{\lambda}^{n}$ as follows
$\delta_{\lambda}^{n}\left[\left(\eta\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}\right]$
$+\left(1-\delta_{\lambda}^{n}\right)\left[\frac{1}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}\right]$
$\leq f\left(x^{n}\right)+g\left(x^{n}\right)-f\left(x^{n+1}\right)-g\left(x^{n+1}\right).$
Whether $\lambda_{n}=0$ or $\lambda_{n}\neq 0$, and with the condition $\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}>\beta_{n}^{2}$, we can obtain
$\delta_{\lambda}^{n}\left(\eta\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{1}{2}\left(\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}-\beta_{n}^{2}\right)\left\|x^{n}-x^{n-1}\right\|_{M}^{{}^{2}}$
$-\frac{1}{2\left(1+\lambda_{n-1}\right)^{2}}\left\|x^{n}-x^{n-1}\right\|_{M}^{{}^{2}}+\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}$
$=\delta_{\lambda}^{n}\left(\eta\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}+\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}$
$+\frac{\left(1-\delta_{\lambda}^{n}\right)}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}\leq f\left(x^{n}\right)+g\left(x^{n}\right)-f\left(x^{n+1}\right)-g\left(x^{n+1}\right).$
Recalling that $\bar{x}^{n-1}-x^{n-1}=\frac{x^{n}-x^{n-1}}{1+\lambda_{n-1}}$ along with $\|\bar{x}^{n-1}-x^{n-1}\|_{M}^{2}=\frac{1}{(1+\lambda_{n-1})^{2}}\|x^{n}-x^{n-1}\|_{M}^{2}$, and noting that
$\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}\geq\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2},$ (4.6)
then we obtain
$\delta_{\lambda}^{n}\left(\eta\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{1}{2}\left(\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}-\beta_{n}^{2}\right)\left\|x^{n}-x^{n-1}\right\|_{M}^{{}^{2}}$
$\leq f\left(x^{n}\right)+g\left(x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}-f\left(x^{n+1}\right)-g\left(x^{n+1}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Reformulating the above statements and from Proposition 1, we arrive at
$\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}\left\|x^{n}-x^{n-1}\right\|^{2}\leq\frac{C_{\lambda}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}\leq f\left(x^{n}\right)+g\left(x^{n}\right)$ (4.7)
$+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}-f\left(x^{n+1}\right)-g\left(x^{n+1}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$ (4.8)
Summing (4.8) from $n_{0}$ to $N$, we further get
$\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}\cdot\sum_{n=n_{0}}^{N}{\left\|x^{n}-x^{n-1}\right\|^{2}}\leq f\left(x^{n_{0}}\right)+g\left(x^{n_{0}}\right)$
$+\frac{1}{2}\left\|\bar{x}^{n_{0}-1}-x^{n_{0}-1}\right\|_{M}^{2}-f\left(x^{N+1}\right)-g\left(x^{N+1}\right)-\frac{1}{2}\left\|\bar{x}^{N}-x^{N}\right\|_{M}^{2}.$
In addition, notice that $f(x)+g(x)$ is proper and bounded below, therefore, $f\left(x^{n}\right)+g\left(x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}$ is non-increasing and bounded below. Consequently, letting $N\rightarrow\infty$, we have the following inequality
$\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}\cdot\sum_{n=n_{0}}^{\infty}{\left\|x^{n}-x^{n-1}\right\|^{2}}\leq f\left(x^{n_{0}}\right)+g\left(x^{n_{0}}\right)+\frac{1}{2}\left\|\bar{x}^{n_{0}-1}-x^{n_{0}-1}\right\|_{M}^{2}$
$-\mathop{\mathrm{inf}}\limits_{N\rightarrow\infty}\left(f\left(x^{N+1}\right)+g\left(x^{N+1}\right)+\frac{1}{2}\left\|\bar{x}^{N}-x^{N}\right\|_{M}^{2}\right)<\infty.$
This leads to (4.3) and the proof is finished. ∎
We now establish the global convergence of the sequence $x^{n}$ generated by Algorithm 1. This analysis employs a perturbation energy function
$A\left(x,\bar{y},y\right)=f\left(x\right)+g\left(x\right)+\frac{1}{2}\left\|\bar{y}-y\right\|_{M}^{2}$ (4.9)
to derive boundedness and convergence guarantees for the iterates.
Theorem 1.
Let $\left\{x^{n}\right\}_{n}$ is generated from $npDCAe_{nls}$ in Algorithm 1 for solving (1.1), assume that $\nabla g_{1}(x)$ is Lipschitz continuous with constant $L_{g_{1}}$, $M$ is a positive definite and linear operator, and $A\left(x,\bar{y},y\right)$ has the KL property, then the following properties hold
  • (i)
    Monotonicity: The sequence $\left\{A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\right\}_{n}$ is monotonically decreasing and converges to some limit: $\lim_{n\rightarrow\infty}A\left(x^{n+1},\bar{x}^{n},x^{n}\right)=\zeta$.
  • (ii)
    Boundedness: The iterates $\left\{x^{n}\right\}_{n}$ is bounded.
  • (iii)
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition.
  • (iv)
    Convergence: The sequence $\left\{x^{n}\right\}_{n}$ is globally convergent with $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$.
Proof.
(i). Setting $C_{1}=\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}$, with (4.8), we obtain
$A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\leq A\left(x^{n},\bar{x}^{n-1}x^{n-1}\right)-C_{1}\left\|x^{n}-x^{n-1}\right\|^{2}.$ (4.10)
This indicates that $\left\{A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\right\}_{n>n_{0}}$ is monotonically decreasing, given the lower boundedness of $f\left(x\right)+g\left(x\right)$. Consequently, we have
$\lim_{n\rightarrow\infty}A\left(x^{n+1},\bar{x}^{n},x^{n}\right)=\zeta.$ (4.11)
(ii). We know that the monotone decrease of $\left\{A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\right\}_{n>n_{0}}$ established in (i) implies
$f\left(x^{n+1}\right)+g\left(x^{n+1}\right)\leq A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\leq f\left(x^{n_{0}+1}\right)+g\left(x^{n_{0}+1}\right)+\frac{1}{2}\left\|\bar{x}^{n_{0}}-x^{n_{0}}\right\|_{M}^{2},\ \forall n>n_{0}$
The level-boundedness of $f\left(x\right)+g\left(x\right)$ together with (4.3) guarantees the boundedness of $\left\{x^{n}\right\}_{n}$.
(iii). Given the boundedness of $\left\{x^{n}\right\}_{n}$ from (ii), there exists at least one cluster point $x^{*}$ with a convergent subsequence $\left\{x^{n_{i}}\right\}$ satisfying
$\lim_{i\rightarrow\infty}x^{n_{i}}=x^{*}\text{ and }\lim_{i\rightarrow\infty}\left\|x^{n_{i}}-x^{n_{i}-1}\right\|\rightarrow 0.$
The momentum update $y^{n_{i}}=\left(x^{n_{i}}+\beta_{n_{i}}\left(x^{n_{i}}-x^{n_{i}-1}\right)\right)$ hence satisfies
$\lim_{i\rightarrow\infty}y^{n_{i}}=x^{*}$ (4.12)
due to the boundedness of $\left\{\beta_{n}\right\}_{n}$ and vanishing step differences (4.3). From the first-order optimality condition of the proximal update (3.2) at $y=\bar{x}^{n_{i}}$, we have
$M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+\nabla f\left(\bar{x}^{n_{i}}\right)+\nabla g_{1}\left(\bar{x}^{n_{i}}\right)\in\partial g_{2}({x}^{n_{i}}).$
We thus arrive at
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla f\left(\bar{x}^{n_{i}}\right)-\nabla g_{1}\left(\bar{x}^{n_{i}}\right)+\nabla(f+g_{1})({x}^{n_{i}})\in-\partial g_{2}({x}^{n_{i}})+\nabla(f+g_{1})({x}^{n_{i}}).$
The left-hand side of the above equation can be written as
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla f({x}^{n_{i}})-\nabla f(\bar{x}^{n_{i}}))+(\nabla g_{1}({x}^{n_{i}})-\nabla g_{1}(\bar{x}^{n_{i}})).$ (4.13)
With $\bar{x}^{n_{i}}-x^{n_{i}}=(x^{n_{i}+1}-x^{n_{i}})/(1+\lambda_{n_{i}})$ and $\bar{x}^{n_{i}}-y^{n_{i}}=\bar{x}^{n_{i}}-x^{n_{i}}-\beta_{n_{i}}(x^{n_{i}}-x^{n_{i}-1})$, the Lipschitz continuity of $\nabla f$ and $\nabla g_{1}$, and (4.3), we conclude that the formulation in (4.13) tends to zero while $n_{i}\rightarrow\infty$. We then obtain
$0\in\lim_{n_{i}\rightarrow\infty}-\partial g_{2}({x}^{n_{i}})+\nabla(f+g_{1})({x}^{n_{i}}).$
Remembering $x^{n_{i}}\rightarrow x^{*}$, with closedness of $\partial g_{2}$ [7, Proposition 16.36] or [26, Theorem 8.6], we have
$0\in-\partial g_{2}\left(x^{*}\right)+\nabla f\left(x^{*}\right)+\nabla g_{1}\left(x^{*}\right),$
confirming that $x^{*}$ satisfies the first-order optimality condition for (1.1).
(iv). The subgradient of the energy function is given by
$\partial A\left(x,\bar{y},y\right)=\left(\begin{array}[]{c}\nabla f\left(x\right)+\nabla g_{1}\left(x\right)-\partial g_{2}(x)\\ M\left(\bar{y}-y\right)\\ -M\left(\bar{y}-y\right)\\ \end{array}\right).$ (4.14)
From the optimality condition
$0\in-\xi^{n}+M\left(\bar{x}^{n}-y^{n}\right)+\nabla f\left(\bar{x}^{n}\right)+\nabla g_{1}\left(\bar{x}^{n}\right),$
we derive the inequality
$\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)\leq\left(L+L_{g_{1}}\right)\left\|\bar{x}^{n}-x^{n}\right\|+\left\|M\left(\bar{x}^{n}-y^{n}\right)\right\|+\left\|M(\bar{x}^{n-1}-x^{n-1})\right\|.$
Setting $\tilde{L}=L+L_{g_{1}}$, with $\bar{x}^{n}-x^{n}=\frac{\lambda_{n}}{1+\lambda_{n}}\left(x^{n+1}-x^{n}\right)$, $\bar{x}^{n-1}-x^{n-1}=\frac{\lambda_{n-1}}{1+\lambda_{n-1}}\left(x^{n}-x^{n-1}\right)$, we obtain the following formulation
$\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)\leq\tilde{L}\left\|x^{n+1}-x^{n}\right\|+\left\|M\left(x^{n+1}-x^{n}\right)\right\|$
$+\beta_{n}\left\|M\left(x^{n}-x^{n-1}\right)\right\|+\left\|M\left(x^{n}-x^{n-1}\right)\right\|.$
Together with the boundedness of $M$ and $\beta_{n}$, we have
$\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)\leq C\left(\left\|x^{n+1}-x^{n}\right\|+\left\|x^{n}-x^{n-1}\right\|\right),$
where $C$ is a constant. With $\lim_{n\rightarrow\infty}\left\|x^{n}-x^{n-1}\right\|\rightarrow 0$ by (4.3), we obtain
$\lim_{n\rightarrow\infty}\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)=0.$ (4.15)
Given the KL property of $A\left(x,\bar{y},y\right)$ with concave function $\psi\left(\cdot\right)$, and from (4.11), we establish
$\left[\psi\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)-\psi\left(A\left(x^{n+1},\bar{x}^{n},x^{n}\right)-\zeta\right)\right]\cdot\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)$
$\geq\psi\prime\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)\cdot\left[A\left(x^{n+1},\bar{x}^{n},x^{n}\right)-A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right]\cdot$
$\mathrm{dist}\left(0,\partial A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right)\geq A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\geq C_{1}\left\|x^{n}-x^{n-1}\right\|^{2}.$
Set $\varPsi\left(x^{n-1},\bar{x}^{n-1},x^{n},\bar{x}^{n},x^{n+1},\zeta\right)=\psi\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)-\psi\left(A\left(x^{n+1},\bar{x}^{n},x^{n}\right)-\zeta\right)$. We further obtain the following statement
$\left\|x^{n}-x^{n-1}\right\|^{2}\leq\tilde{C}\cdot\varPsi\left(x^{n-1},\bar{x}^{n-1}x^{n},\bar{x}^{n},x^{n+1},\zeta\right)\cdot\left(\left\|x^{n+1}-x^{n}\right\|+\left\|x^{n}-x^{n-1}\right\|\right),$
where $\tilde{C}=\frac{C}{C_{1}}$. From the basic inequality $a^{2}\leq cd\Longrightarrow a\leq c+\frac{d}{4}$, we obtain
$\frac{1}{2}\left\|x^{n}-x^{n-1}\right\|$ $\leq\tilde{C}\cdot\left[\psi\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)-\psi\left(A\left(x^{n+1},\bar{x}^{n},x^{n}\right)-\zeta\right)\right]$ (4.16)
$+\frac{1}{4}\left(\left\|x^{n+1}-x^{n}\right\|-\left\|x^{n}-x^{n-1}\right\|\right).$ (4.17)
Summing both sides of the inequality, we obtain
$\frac{1}{2}\sum_{n=n_{0}}^{N}\left\|x^{n}-x^{n-1}\right\|$ $\leq\tilde{C}\left[\psi\left(A\left(x^{n_{0}},\bar{x}^{n_{0}-1},x^{n_{0}-1}\right)-\zeta\right)-\psi\left(A\left(x^{N+1},\bar{x}^{N},x^{N}\right)-\zeta\right)\right]$
$+\frac{1}{4}\left(\left\|x^{N+1}-x^{N}\right\|-\left\|x^{n_{0}}-x^{n_{0}-1}\right\|\right).$
Letting $N\rightarrow\infty$, we know that
$\lim_{N\rightarrow\infty}\psi\left(A\left(x^{N+1},\bar{x}^{N},x^{N}\right)-\zeta\right)=\psi\left(0\right)=0.$
Therefore, with (4.3), we obtain the desired statement
$\sum_{n=n_{0}}^{\infty}{\left\|x^{n}-x^{n-1}\right\|\leq 2\tilde{C}\cdot\psi\left(A\left(x^{n_{0}},\bar{x}^{n_{0}-1},x^{n_{0}-1}\right)-\zeta\right)}<\infty.$
Finally, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$. ∎
We now establish convergence properties for Algorithm 2. While Algorithm 2 differs from Algorithm 1 only in its DC decomposition strategy, their convergence proofs share identical methodology. We present three parallel conclusions for Algorithm 2 for completeness.
4.2 Convergence analysis of Algorithm 2
In contrast to Algorithm 1, Algorithm 2 requires the function $f$ to be convex. The following Lemma 3, Lemma 4 and Theorem 2 are obtained on this basis, and we will not write out one by one later on. Following the analytical framework established for Algorithm 2, we initially demonstrate that Lemma 3 serves to rigorously characterize the quasi-descent property intrinsic to the energy functional governing the system.
Lemma 3.
Let $\bar{x}^{n}$ is generated by (3.10), then the following inequality holds
$E\left(\bar{x}^{n}\right)\leq E\left(x^{n}\right)+\frac{L\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{L}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$ (4.18)
Proof.
Firstly, from (3.10), we know that $\bar{x}^{n}$ is the minimum point of the strongly convex function
$\nabla f\left(y^{n}\right)-\xi^{n},y\right>+\frac{L}{2}\left\|y-y^{n}\right\|_{M}^{2}+g_{1}\left(y\right).$
This implies the optimality inequality
$\nabla f\left(y^{n}\right)-\xi^{n},\bar{x}^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)\leq\left<\nabla f\left(y^{n}\right)-\xi^{n},x^{n}\right>+\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
$+g_{1}\left(x^{n}\right)-\frac{L}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Using the subgradient property $\xi^{n}\in\partial g_{2}\left(x^{n}\right)$, then we obtain
$g_{2}\left(\bar{x}^{n}\right)\geq g_{2}\left(x^{n}\right)+\left<\xi^{n},\bar{x}^{n}-x^{n}\right>.$
Combining these results yields
$\nabla f\left(y^{n}\right),\bar{x}^{n}-x^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(\bar{x}^{n}\right)\leq\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
$+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\frac{L}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Since $f(x)$ is convex and $\nabla f(x)$ is Lipschitz continuous, we have the following two inequalities
$f\left(\bar{x}^{n}\right)\leq f\left(y^{n}\right)+\left<\nabla f\left(y^{n}\right),\bar{x}^{n}-y^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|^{2},$
$f\left(x^{n}\right)\geq f\left(y^{n}\right)+\left<\nabla f\left(y^{n}\right),x^{n}-y^{n}\right>.$
Sum the above two inequalities, and obtain
$f\left(\bar{x}^{n}\right)-g_{2}\left(\bar{x}^{n}\right)+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)\leq\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(x^{n}\right).$
$-\frac{L}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-g_{2}\left(x^{n}\right)+f\left(x^{n}\right)+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|^{2}.$
Applying the matrix dominance $M\succeq I$ and $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, we derive (4.18). ∎
In the subsequent analysis, we establish the global subsequential convergence property for Algorithm 2. To provide a basis for this study, we first present the following preparatory Lemma 4, whose proof procedure is essentially similar to that of Algorithm 1, and we omit it here.
Lemma 4.
Let $\left\{x^{n}\right\}_{n}$ is generated from $pDCAe_{\text{nls}}$ in Algorithm 2 for solving (1.1), then we have
$\sum_{n=1}^{\infty}{\left\|x^{n+1}-x^{n}\right\|^{2}<\infty}.$ (4.19)
To conclude this subsection, we establish the global convergence of Algorithm 2. For consistency, we adopt identical conventions as those defined for Algorithm 1, given the essential similarity in their analytical frameworks. The detailed theoretical proof of this lemma parallels the methodology employed for Algorithm 1.
Theorem 2.
[…]
$A\left(x,\bar{y},y\right)=f\left(x\right)+g\left(x\right)+\frac{L}{2}\left\|\bar{y}-y\right\|_{M}^{2}.$
Assume that $\nabla g_{1}(x)$ is Lipschitz continuous with constant $L_{g_{1}}$, $M\succeq I$ being a positive definite and linear operator, and $A\left(x,\bar{y},y\right)$ has the KL property, then the following properties hold
  • (i)
    Monotonicity: The sequence $\left\{A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\right\}_{n}$ is monotonically decreasing and converges to some limit: $\lim_{n\rightarrow\infty}A\left(x^{n+1},\bar{x}^{n},x^{n}\right)=\zeta$.
  • (ii)
    Boundedness: The iterates $\left\{x^{n}\right\}_{n}$ is bounded.
  • (iii)
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition.
  • (iv)
    Convergence: The sequence $\left\{x^{n}\right\}_{n}$ is globally convergent with $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$.
4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
Throughout this paper, suppose that $g_{1}(x)$ is proper, lower semicontinuous and convex as follows
$g_{1}\left(x\right)\geq g_{1}\left(\bar{x}\right)+\langle x-\bar{x},\tilde{w}\rangle+\frac{\mu}{2}\left\|x-\bar{x}\right\|^{2},\quad\mu\geq 0,$ (4.20)
for any $\bar{x}\in\operatorname{dom}g_{1}$ and $\tilde{w}\in\partial g_{1}(\bar{x})$ and $\mu$ can be zero while $g_{1}$ is not strongly convex. We propose a more general case under the following Assumption 1. It can cover the widely used case $g_{1}=\|x\|_{1}$ along with $\nabla g_{2}(x)$ being Lipschitz continuous with constant $L_{g_{2}}$, which includes some important regularization such as SCAD as discussed in [1].
Assumption 1.
We assume the proper, lower semicontinuous and convex function $g_{1}$ has the following property, for any $\bar{x}\in\operatorname{dom}g_{1}$, any sufficiently small $\varepsilon>0$, and $x\in B_{\varepsilon}(\bar{x})\cap\operatorname{dom}g_{1}$, there exits a subgradient $\bar{w}\in\partial g_{1}(\bar{x})$ and $\sigma>0$ such that
$g_{1}\left(x\right)\leq g_{1}\left(\bar{x}\right)+\langle x-\bar{x},\bar{w}\rangle+\frac{\sigma}{2}\left\|x-\bar{x}\right\|^{2}.$ (4.21)
In the following proposition, we rigorously demonstrate that the $l_{1}$-norm satisfies Assumption 1. This analysis confirms the theoretical soundness of our proposed algorithmic framework and verifies that the required assumption is readily satisfiable in practical implementations.
Proposition 3.
The anisotropic $l_{1}$-norm in $\mathbb{R}^{n}$
$h\left(x\right)=\left\|x\right\|_{1}=\left|x_{1}\right|+\left|x_{2}\right|+\ldots+\left|x_{n}\right|$
satisfies the Assumption 1.
Proof.
Let us first consider the case $h(x)=|x|$ with $x\in\mathbb{R}$. It is known that
$\partial h(x)$ =$=\begin{cases}1&x>0\\ -1&x<0\\ [-1,1]&x=0\end{cases}.$
It can be checked that $\nabla h(\bar{x})=\mathrm{sign}(\bar{x})$ for any $\bar{x}\neq 0$. There thus exists small $\varepsilon>0$, for any $|x-\bar{x}|<\varepsilon$, such that
$|x|\leq|\bar{x}|+\langle x-\bar{x},\mathrm{sign}(\bar{x})\rangle+\frac{\sigma}{2}\|x-\bar{x}\|^{2},$
for any $\sigma>0$. Moreover, if $\bar{x}=0$, for small $\varepsilon>0$ and $|x-\bar{x}|<\varepsilon$, we can choose $\bar{w}=\mathrm{sign}(x)\in\partial h(\bar{x})|_{\bar{x}=0}$. One can readily verify that
$|x|\leq\langle x,\mathrm{sign}(x)\rangle+\frac{\sigma}{2}\|x-\bar{x}\|^{2},$
for any $\sigma>0$. For the case $x\in\mathbb{R}^{n}$, the subdifferential of $h(x)$ is the set of all subgradients that can be generated as follows
$\partial h(x)$ =$=\big{\{}g\mid\|g\|_{\infty}\leq 1,\ g^{T}x=\|x\|_{1}\big{\}}.$
For $\bar{x}=\left(\bar{x}_{1},\bar{x}_{2},\ldots,\bar{x}_{n}\right)$, $x\in B_{\varepsilon}(\bar{x})$ with small enough $\varepsilon$, we can choose
$\bar{w}=\left(g_{1},g_{2},\ldots,g_{n}\right),\quad\begin{cases}g_{i}=\nabla|\bar{x}_{i}|=\mathrm{sign}(\bar{x}_{i}),\ &\text{if}\ \bar{x}_{i}\neq 0,\\ g_{i}=\mathrm{sign}(x_{i}),\ &\text{if}\ \bar{x}_{i}=0.\end{cases}$
It can be verified that
$\sum_{i=1}^{n}{\left|x_{i}\right|\leq\sum_{i=1}^{n}{\left|\bar{x}_{i}\right|}+\sum_{i=1}^{n}{\left(x_{i}-\bar{x}_{i}\right)\cdot g_{i}}+\frac{\sigma}{2}\sum_{i=1}^{n}{\left(x_{i}-\bar{x}_{i}\right)^{2}}},$ (4.22)
for any $\sigma>0$. We thus conclude that the anisotropic $l_{1}$-norm satisfies (4.21). ∎
Now we introduce the following perturbation energy function
$H\left(x,w,\bar{y},y\right)=f\left(x\right)+\left<x,w\right>-g_{1}^{*}\left(w\right)-g_{2}\left(x\right)+\frac{\sigma}{2}\left\|x-\bar{y}\right\|^{2}+\frac{1}{2}\left\|\bar{y}-y\right\|_{M}^{2},$ (4.23)
for proving the global convergence of the Algorithm 1 for the case $g_{1}$ being non-smooth.
Theorem 3.
Let $\left\{x^{n}\right\}_{n}$ is generated from $npDCAe_{nls}$ in Algorithm 1 for solving (1.1). Suppose that $\left\{x^{n}\right\}_{n\in N}$ has a cluster $x^{*}$ and $H\left(x,w,\bar{y},y\right)$ is a KL function, $N_{0}=\max(n_{0},n_{1})$, then the following statements hold
  • (i)
    Monotonicity and boundedness: The sequence $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n\geq N_{0}}$ is decreasing monotonically and converges to some limits: $\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}$. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
  • (ii)
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimality condition.
  • (iii)
    Convergence: The sequence $\left\{x^{n}\right\}_{n}$ is globally convergent with $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$.
Proof.
(i). Using the conjugate property $\bar{x}^{n},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)=g_{1}\left(\bar{x}^{n}\right)$, we derive
$x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)=\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>+g_{1}\left(\bar{x}^{n}\right).$ (4.24)
For convenience, we define an auxiliary and intermediate function
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right):=f\left(x^{n+1}\right)+\left<x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)-g_{2}\left(x^{n+1}\right).$ (4.25)
First with (4.25), then by (4.20) with $\bar{x}=\bar{x}^{n}$ and $x=x^{n+1}$, for any $\bar{w}^{n}\in\partial g_{1}(\bar{x}^{n})$, we have
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $=f\left(x^{n+1}\right)+\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(x^{n+1}\right)$
$\leq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right)-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2},$
where the equality comes from (4.24). Using the energy decreasing (4.1), together with the line search condition (3.9), for the case 1 of (3.6), we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.26)
For the case 2 of (3.6), while the line search does not succeed with $N_{\max}$ searches, we assume $\lambda_{n}=0$ along with $\nu_{n}$ set to 0 and hence $x^{n+1}=\bar{x}^{n}$ as in (4.1), i.e.,
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\frac{1}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.27)
We now combine the inequality (4.26) and (4.27) with $\delta_{\lambda}^{n}$ as follows
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\delta_{\lambda}^{n}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}$ (4.28)
$-\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.29)
For sufficiently large $N_{0}$, while $n>N_{0}$, we conclude $x^{n}\in B_{\varepsilon}(\bar{x}^{n-1})$ for small enough $\varepsilon$ by the square summability of $\{x^{n}\}_{n}$ as in (4.3) with $\bar{x}^{n-1}-x^{n-1}=(x^{n}-x^{n-1})/(1+\lambda_{n-1})$. Therefore, applying (4.21) as in the Assumption 1 with $x=x^{n}$, $\bar{x}=\bar{x}^{n-1}$, there exists $\bar{w}=\bar{w}^{n-1}$ and $\bar{w}^{n-1}\in\partial g_{1}\left(\bar{x}^{n-1}\right)$, such that
$g_{1}(x^{n})$ $\leq g_{1}(\bar{x}^{n-1})+\langle x^{n}-\bar{x}^{n-1},\bar{w}^{n-1}\rangle+\frac{\sigma}{2}\|x^{n}-\bar{x}^{n-1}\|^{2}$
$=\langle x^{n},\bar{w}^{n-1}\rangle-g_{1}^{*}(\bar{w}^{n-1})+\frac{\sigma}{2}\|x^{n}-\bar{x}^{n-1}\|^{2}.$ (4.30)
Substituting $g_{1}(x^{n})$ as in (4.30) into (4.29), we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)-\delta_{\lambda}^{n}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}.$
With $(x^{n}-x^{n-1})/(1+\lambda_{n-1})=\bar{x}^{n-1}-x^{n-1}$ and $\beta_{n}^{2}<\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}$, still with (4.6), reformulating the above inequality, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)$
$+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}-\delta_{n}^{\lambda}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}$
$+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{1}{2}\left(\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}-\beta_{n}^{2}\right)\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$
If taking $\eta$ be larger, whether $\lambda_{n}=0$ or $\lambda_{n}>0$, we have
$\delta_{n}^{\lambda}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}\geq 0,$ (4.31)
since the left hand side of (4.31) is 0 for the case $\delta_{n}^{\lambda}=0$, i.e., $\lambda_{n}=0$ and $x^{n+1}=\bar{x}^{n}$. Then from Proposition 1, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{C_{\lambda}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}$
$\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}.$
Reformulating the above inequality, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}\left\|x^{n}-x^{n-1}\right\|^{2}$
$\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}.$
Therefore, for $n>N_{0}$, we obtain
$H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\leq H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-C_{1}\left\|x^{n}-x^{n-1}\right\|^{2},$ (4.32)
which means $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is decreasing monotonically. Furthermore, as we can see that $\bar{w}^{n}\in\partial g_{1}\left(\bar{x}^{n}\right)$, from Assumption 1, thus we obtain
$f\left(x^{n+1}\right)+\left<x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$=f\left(x^{n+1}\right)+g_{1}\left(\bar{x}^{n}\right)+\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$\geq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$\geq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right).$ (4.33)
We see that $f(x)+g(x)$ is bounded below, which implies
$\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}.$ (4.34)
On the other hand, since $f(x)+g(x)$ is proper, closed, and level-bounded, we conclude that the iteration sequence $\{x^{n}\}_{n}$ is bounded with (4.33) and monotone decreasing of $H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)$.
(ii). Since there exists a cluster point $x^{*}$ with a convergent subsequence $\left\{x^{n_{i}}\right\}_{n_{i}}$ satisfying
$\lim_{i\rightarrow\infty}x^{n_{i}}=x^{*}\text{ and }\lim_{i\rightarrow\infty}\left\|x^{n_{i}}-x^{n_{i}-1}\right\|\rightarrow 0.$
The momentum update $y^{n_{i}}=\left(x^{n_{i}}+\beta^{n_{i}}\left(x^{n_{i}}-x^{n_{i}-1}\right)\right)$ consequently satisfies $\lim_{i\rightarrow\infty}y^{n_{i}}=x^{*}$, due to the boundedness of $\left\{\beta_{n}\right\}_{n}$ and vanishing step differences in (4.3). From the first-order optimality condition (3.2) at $y=\bar{x}^{n_{i}}$, we have
$0\in-\xi^{n_{i}}+M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+\nabla f\left(\bar{x}^{n_{i}}\right)+\partial g_{1}\left(\bar{x}^{n_{i}}\right).$
With $\xi^{n_{i}}=\nabla g_{2}(x^{n_{i}})$, we thus obtain
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)+\nabla g_{2}\left(x^{n_{i}}\right)\in\nabla f\left(\bar{x}^{n_{i}}\right)+\partial g_{1}\left(\bar{x}^{n_{i}}\right)-\nabla g_{2}(\bar{x}^{n_{i}}).$
The left-hand side of the above equation can be written as
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla g_{2}\left(x^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)).$ (4.35)
Furthermore, the cumulative step bound (4.3) implies
$\lim_{i\rightarrow\infty}\left\|\bar{x}^{n_{i}}-x^{n_{i}}\right\|\rightarrow 0\text{ and }\lim_{i\rightarrow\infty}\bar{x}^{n_{i}}=x^{*}.$ (4.36)
With $\bar{x}^{n_{i}}-x^{n_{i}}=(x^{n_{i}+1}-x^{n_{i}})/(1+\lambda_{n_{i}})$ and $\bar{x}^{n_{i}}-y^{n_{i}}=\bar{x}^{n_{i}}-x^{n_{i}}-\beta_{n_{i}}(x^{n_{i}}-x^{n_{i}-1})$, the Lipschitz continuity of $\nabla g_{2}$, and (4.36), we conclude that the formulation in (4.35) tends to zero while $n_{i}\rightarrow\infty$. We then obtain
$0\in\lim_{n_{i}\rightarrow\infty}-\partial g_{1}(\bar{x}^{n_{i}})+\nabla(f+g_{2})(\bar{x}^{n_{i}}).$
With closedness of $\partial g_{1}$ [7, Proposition 16.36] or [26, Theorem 8.6] and (4.36), we have
$0\in-\partial g_{1}\left(x^{*}\right)+\nabla f\left(x^{*}\right)+\nabla g_{2}\left(x^{*}\right),$
confirming $x^{*}$ satisfies the first-order optimality condition for (1.1).
(iii). The subgradient of the energy function is given by
$\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)=\left(\begin{array}[]{c}\nabla f\left(x^{n}\right)+\bar{w}^{n-1}-\nabla g_{2}\left(x^{n}\right)+\sigma\left(x^{n}-\bar{x}^{n-1}\right)\\ x^{n}-\partial g_{1}^{*}\left(\bar{w}^{n-1}\right)\\ -\sigma\left(x^{n}-\bar{x}^{n-1}\right)+M\left(\bar{x}^{n-1}-x^{n-1}\right)\\ -M\left(\bar{x}^{n-1}-x^{n-1}\right)\\ \end{array}\right).$ (4.37)
From first-order optimality condition, we have
$0\in\nabla f\left(\bar{x}^{n-1}\right)+\partial g_{1}\left(\bar{x}^{n-1}\right)-\nabla g\left(x^{n-1}\right)+M\left(\bar{x}^{n-1}-y^{n-1}\right).$
Since $\bar{\omega}^{n-1}\in\partial g_{1}\left(\bar{x}^{n-1}\right)$ and $\bar{x}^{n-1}\in\partial g_{1}^{*}\left(\bar{\omega}^{n-1}\right)$, with conditions $\bar{x}^{n-1}-x^{n-1}=\frac{x^{n}-x^{n-1}}{1+\lambda_{n-1}}$ and $x^{n}-\bar{x}^{n-1}=\lambda_{n-1}\left(\bar{x}^{n-1}-x^{n-1}\right)$, then reformulating the statement, we thus have
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\leq L\left\|x^{n}-\bar{x}^{n-1}\right\|+L_{g_{2}}\left\|x^{n}-x^{n-1}\right\|$
$+\left(2\sigma+1\right)\left\|x^{n}-\bar{x}^{n-1}\right\|+3B_{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|+B_{1}\left\|x^{n-1}-x^{n-2}\right\|\quad\left(*\right),$
where $B_{1}$ is the maximum singular value of $M$. Adjusting the expression above yields
$\left(*\right)\leq B_{2}\left\|x^{n}-x^{n-1}\right\|+B_{1}\left\|x^{n-1}-x^{n-2}\right\|\leq B\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right),$
where $B_{2}=(L+2\sigma+1)\cdot\lambda_{\max}+3B_{1}+L_{g_{2}}$ and $B=\max(B_{1},B_{2})$ are constants. Then we obtain the statement
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\leq B\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right).$ (4.38)
From (4.3), we know that $\lim_{n\rightarrow\infty}\left\|x^{n}-x^{n-1}\right\|\rightarrow 0$, we thus get
$\lim_{n\rightarrow\infty}\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)=0.$
Given the KL property of $H\left(x,w,\bar{y},y\right)$ with concave function $\psi\left(\cdot\right)$, and from (4.11), we establish
$\left[\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)-\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\zeta\right)\right]\cdot$
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\geq\psi\prime\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)\cdot$
$\left[H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right]\cdot\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)$
$\geq H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\geq C_{1}\left\|x^{n}-x^{n-1}\right\|^{2}.$
Then we set
$\varPsi\left(x^{n-1},\bar{x}^{n-1},\bar{w}^{n-1},x^{n},\bar{x}^{n},\bar{w}^{n},x^{n+1},\bar{\zeta}\right)=\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-$
$\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right).$
We further obtain the following statement
$\left\|x^{n}-x^{n-1}\right\|^{2}$ $\leq\tilde{B}\cdot\varPsi\left(x^{n-1},\bar{x}^{n-1},\bar{w}^{n-1},x^{n},\bar{x}^{n},\bar{w}^{n},x^{n+1},\bar{\zeta}\right)$
$\cdot\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right),$
where $\tilde{B}=\frac{B}{C_{1}}$. From the basic inequality $a^{2}\leq cd\Longrightarrow a\leq c+\frac{d}{4}$, we obtain
$\frac{1}{2}\left\|x^{n}-x^{n-1}\right\|$ $\leq\tilde{B}\cdot\left[\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right)\right]$
$+\frac{1}{4}\left(\left\|x^{n-1}-x^{n-2}\right\|-\left\|x^{n}-x^{n-1}\right\|\right).$
Summing both sides of the inequality, we obtain
$\frac{1}{2}\sum_{n=N_{0}}^{N}{\left\|x^{n}-x^{n-1}\right\|}\leq\frac{1}{4}\left(\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|-\left\|x^{N}-x^{N-1}\right\|\right)$
$+\tilde{B}\left[\psi\left(H\left(x^{N_{0}},\bar{w}^{N_{0}-1},\bar{x}^{N_{0}-1},x^{N_{0}-1}\right)-\bar{\zeta}\right)-\psi\left(H\left(x^{N+1},\bar{w}^{N},\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)\right].$
Letting $N\rightarrow\infty$, we know that
$\lim_{N\rightarrow\infty}\psi\left(H\left(x^{N+1},\bar{w}^{N},\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)=\psi\left(0\right)=0.$
Therefore, from (4.3), we obtain the desired statement
$\sum_{n=N_{0}}^{\infty}\|x^{n}-x^{n-1}\|\leq$ $2\tilde{B}\cdot\psi(H\left(x^{N_{0}},\bar{w}^{N_{0}-1},\bar{x}^{N_{0}-1}x^{N_{0}-1}\right)-\bar{\zeta})+\frac{1}{2}\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|<\infty.$
Consequently, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$ and the global convergence of$\{x^{n}\}_{n}$. ∎
Below we give the conclusion of this particular calculus in the framework of Algorithm 2 and omit a similar proof for compactness.
Theorem 4.
Assume $\left\{x^{n}\right\}_{n}$ is generated from $pDCAe_{nls}$ in Algorithm 2 for solving (1.1). Define the perturbation energy function
$H\left(x,w,\bar{y},y\right)=f\left(x\right)+\left<x,w\right>-g_{1}^{*}\left(w\right)-g_{2}\left(x\right)+\frac{L\sigma}{2}\left\|x-\bar{y}\right\|^{2}+\frac{L}{2}\left\|\bar{y}-y\right\|_{M}^{2}.$
Suppose that $\left\{x^{n}\right\}_{n\in N}$ has a cluster $x^{*}$ and $H\left(x,w,\bar{y},y\right)$ is a KL function, then the following statements hold
  • (i)
    Monotonicity: The sequence $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is monotonically decreasing and converges to some limit: $\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}$.
  • (ii)
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
  • (iii)
    Convergence: The sequence $\left\{x^{n}\right\}_{n}$ is globally convergent with $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$.
Remark 3.
As for the framework of Algorithm 2, we can naturally obtain global convergence. Noting that the difference between Algorithm 1 and Algorithm 2 is the difference of (4.3) and (4.19). In the above proof, we have only used (4.3) after the (4.21); therefore, in the framework of Algorithm 2, the subsequent proofs are not quite different from the proofs for Algorithm 1.
4.4 Convergence rate of the Algorithm 1
Our analysis requires explicit characterization of the Kurdyka-Łojasiewicz (KL) exponent associated with the auxiliary function $A(x,\bar{y},y)$ to quantify local convergence rates. The subsequent theorem, which adapts well-established convergence rate theory (cf. [5, Theorem 2], [2, Lemma 1]), formalizes this relationship.
Theorem 5 (Local convergence rate).
Under the assumptions of Theorem 1, let $\{x^{n}\}_{n}$ be a sequence generated by Algorithm 1 with a convergent subsequence $x^{n_{i}}\to x^{*}$. If $A(x,\bar{y},y)$ is a KL function with $\phi(s)=cs^{1-\theta}$ in the KL inequality $(\theta\in[0,1),c>0)$, then
  • (i)
    When $\theta=0$, $\exists\ m_{0}>0$ such that
    $x^{n}=x^{*}\quad\forall\ n>m_{0}.$
  • (ii)
    For $\theta\in(0,\frac{1}{2}]$, $\exists\ p_{1}>0$, $\eta\in(0,1)$, and $m_{1}>0$ such that
    $\|x^{n}-x^{*}\|\leq p_{1}\eta^{n}\quad\forall\ n>m_{1}.$
  • (iii)
    For $\theta\in(\frac{1}{2},1)$, $\exists\ p_{2}>0$ and $m_{2}>0$ such that
    $\|x^{n}-x^{*}\|\leq p_{2}n^{-\frac{1-\theta}{2\theta-1}}\quad\forall\ n>m_{2}.$
5 Numerical experiments
In this section, we conduct two representative numerical experiments to validate our algorithm’s advantages in solving non-convex optimization problems
  • SCAD-Regularized Least Squares: Implemented on an AMD Ryzen 7 8845H system (3.80 GHz CPU with Radeon 780M iGPU).
  • Nonlocal Ginzburg-Landau Segmentation: Executed on NVIDIA RTX 3080 Ti GPU system (16GB VRAM).
Detailed implementation protocols and comparative results will be systematically presented in subsequent subsections. For all tables, whence the iteration number exceeds the maximum iterations, we use notation — instead of recording the CPU time.
Remark 4.
Before starting this subsection, we show that the minimization functions of the following two problems do have the KL property. The KL property of the original SCAD regularization (5.1) follows directly from [29, Example 4.3]. For the modified SCAD regularization in (5.2), since each $s_{M,i}(x_{i})$ is a piecewise quadratic function, the KL property of $E(x)$ can be derived via the methodology in [20, Section 5.2]. Combining theoretical frameworks from [20, 22, 4], we confirm that both $A(x,\bar{y},y)$ and $H(x,w,\bar{y},y)$ are KL functions, thereby establishing global convergence. Finally, the discrete graphic Ginzburg-Landau functional, being a polynomial (hence semialgebraic) function, automatically satisfies the KL property [4, Section 2.2].
5.1 Least squares problems with modified SCAD regularizer
We analyze a smoothed variant of the SCAD (Smoothly Clipped Absolute Deviation) regularizer through DC decomposition. Following the framework in [1, Section 6.1] and [29], the SCAD penalty admits
[…]
where $\theta>1$ controls concavity and $\mu>0$ regulates sparsity. The second term $\tilde{S}(x)=\mu\|x\|_{1}-S(x)=\sum_{i}^{k}\tilde{s}_{i}(x_{i})$ decomposes into component functions $\tilde{s}_{i}(x_{i})$ with piecewise structure
$\tilde{s}_{i}(x_{i})=\left\{\begin{array}[]{ll}0&|x_{i}|\leq\mu\\ \frac{(|x_{i}|-\mu)^{2}}{2(\theta-1)\mu}&\mu<|x_{i}|<\theta\\ \mu|x_{i}|-\frac{\mu^{2}(\theta+1)}{2}&|x_{i}|\geq\theta\mu\end{array},\right.\ \ \nabla_{i}\tilde{S}_{i}(x_{i})=\text{sign}(x_{i})\dfrac{[\min\{\theta\mu,|x_{i}|\}-\mu]_{+}}{(\theta-1)}.$
We formulate the $l_{1}$ norm regularized least squares problem with the SCAD regularizer as follows
$\min_{x\in\mathbb{R}^{k}}E(x)=\frac{1}{2}\|Ax-b\|^{2}+S(x).$ (5.1)
To ensure differentiability, we replace the $l_{1}$ norm with the Huber function $\mathcal{H}_{\alpha}(x):=\sum_{i}\mathcal{H}(x_{i},\alpha)$
$\mathcal{H}(x_{i},\alpha)=\left\{\begin{array}[]{ll}\frac{|x_{i}|^{2}}{2\alpha}&|x_{i}|\leq\alpha\\ |x_{i}|-\frac{\alpha}{2}&|x_{i}|>\alpha\end{array},\right.$
where $\alpha=0.5\mu$ balances robustness. The modified SCAD-regularized least squares problem becomes
$\min_{x\in\mathbb{R}^{k}}E(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\mathcal{H}_{\alpha}(x)-\tilde{S}(x),$ (5.2)
with $A\in\mathbb{R}^{m\times k}$, $b\in\mathbb{R}^{m}$. The composite penalty $S_{M}(x):=\mu\mathcal{H}_{\alpha}(x)-\tilde{S}(x):=\sum_{i=1}^{k}s_{M,i}(x_{i})$, where $p_{M,i}(x_{i})$ exhibits piecewise behavior
$\frac{s_{M,i}(x_{i})}{\mu}=\left\{\begin{array}[]{ll}|x_{i}|^{2}/(2\alpha)&|x_{i}|\leq\alpha\\ |x_{i}|-{\alpha}/{2}&\alpha<|x_{i}|\leq\mu\\ |x_{i}|-{\alpha}/{2}-(|x_{i}|-\mu)^{2}/(2(\theta-1)\mu)&\mu<|x_{i}|<\theta\mu\\ (\mu(\theta+1)-\alpha)/{2}&|x_{i}|\geq\theta\mu\end{array}.\right.$ (5.3)
We validate on LIBSVM datasets [10] converted to MATLAB format for binary classification. Initialized at $x^{0}=0$, convergence is triggered when
$\frac{\|x^{n}-x^{n-1}\|}{\max\{1,\|x^{n}\|\}}<\epsilon\text{, where }\epsilon=10^{-i}\text{, for }i=4,5,\ldots,9.$
We will compute the original SCAD (5.1) and the modified SCAD (5.2) for comparisons. From [28], we see that the Huber-SCAD model (5.2) can obtain nearly the same sparsity as the original SCAD model. We refer to [21] for other types of DC-type regularization for signal and image processing.
[…]
Remark 5.
It is worth noting that in solving (5.1) and (5.2), Algorithm 1 and Algorithm 2 become equivalent when appropriate preconditioners are selected. Take (5.1) as an example, when we set $f(x)$, $g_{1}(x)$ and $g_{2}(x)$ as follows:
[…]
Choosing $M=\lambda_{A^{T}A}I-A^{T}A$, we can see that the two algorithms have the same iteration.
  • pDCAe: For this algorithm, we set $L=\lambda_{A^{T}A}$, where $\lambda_{A^{T}A}$ represents the largest eigenvalue of the matrix $A^{T}A$, and choose the extrapolation parameter $\left\{\beta_{n}\right\}$ as FISTA, and perform both fixed restart (with $\bar{T}=200$) and adaptive restart strategy as described above (3.7). One can refer to [29, Section 5.1] for more details.
  • pDCAe(no-restart): This special version of the algorithm pDCAe does not perform the fixed restart or the adaptive restart strategy. The extrapolation parameter of this algorithm has been chosen as FISTA. This comparison aims to show the importance of the restart of the extrapolation parameter.
[…]
    $E(x)=\mu\mathcal{H}_{\alpha}(x)+\frac{\lambda_{A^{T}A}}{2}\|x\|^{2}-\left(\frac{\lambda_{A^{T}A}}{2}\|x\|^{2}+\tilde{S}(x)-\frac{1}{2}\|Ax-b\|^{2}\right),$ (5.4)
    where $\lambda_{A^{T}A}$ represents the maximum eigenvalue of the matrix $A^{T}A$.
[…]
Table 2: Solving (5.1) on random instances
[…]
Precision Iter CPU time (s)
Algorithm
$10^{-4}$ 202 40136 1156 183 93 0.8 126.4 3.9 1.7 0.7
$10^{-5}$ 1002 130844 20124 4162 664 3.4 401.8 64.5 45.7 5.0
$10^{-6}$ 8402 Max 167641 25528 2570 27.2 504.1 200.4 18.1
$10^{-7}$ 39602 Max Max 190425 7839 130.6 1688.1 54.3
$10^{-8}$ 136402 Max Max Max 10335 442.4 72.7
$10^{-9}$ 456002 Max Max Max 16061 1588.2 117.3
Table 3: Solving (5.2) on random instances
[…]
Precision Iter CPU time (s)
Algorithm
$10^{-4}$ 202 40967 1216 193 85 0.8 130.6 4.0 1.8 0.8
$10^{-5}$ 1002 131326 20644 4375 676 3.4 422.1 67.8 47.8 5.4
$10^{-6}$ 8202 Max 168123 27154 2905 27.6 528.4 257.1 21.1
$10^{-7}$ 39002 Max Max 192202 10118 127.6 1793.4 76.3
$10^{-8}$ 135802 Max Max Max 11745 428.8 93.1
$10^{-9}$ 451002 Max Max Max 14298 1458.7 113.2
(a) $l_{1}$-convergence-rate
(b) Huber-convergence-rate
Figure 1: The convergence rate of two algorithms for solving (5.1) and (5.2).
[…]
5.2 Graphic Ginzburg-Landau model for image segmentation
In this subsection, we focus on a segmentation problem using a nonlocal Ginzburg-Landau model with prior guidance, formulated as [8]
$\min_{x\in\mathbb{R}^{N}}E(x)=\sum_{i,j}\frac{\tau}{2}w_{ij}(x(i)-x(j))^{2}+\frac{1}{\tau}\mathbb{W}(x)+\frac{\gamma}{2}\sum_{i}\Lambda(i)(x(i)-y(i))^{2},$ (5.5)
where $x$ denotes segmentation labels, $w_{ij}=K(i,j)\cdot N(i,j)$ encodes nonlocal interactions via Gaussian similarity $K(i,j)=\exp\left(-\|P_{i}-P_{j}\|_{2}^{2}/\kappa^{2}\right)$ and binary adjacency $N(i,j)$, $\mathbb{W}(x)=\frac{1}{4}\sum_{i=1}^{N}(x(i)^{2}-1)^{2}$ enforces binary constraints, the diagonal matrix $\Lambda$ and the vector $y$ incorporate prior labels [28]. We choose a $25\times 25$ square box surrounding each pixel for the nonlocal interactions. Parameters $\tau=\gamma=10$ balance regularization strength.
The above energy function $E(x)$ is divided into two components
$f(x)=\frac{\gamma}{2}\sum_{i}\Lambda(i)(x(i)-y(i))^{2},\ g_{1}(x)=\sum_{ij}\frac{\tau}{2}w_{ij}(x(i)-x(j))^{2}\text{ and }g_{2}(x)=\frac{1}{\tau}\mathbb{W}(x).$
For comparison, we implement the following algorithms
[…]
The first criterion “DICE Bound” in Table 4 relies on the segmentation coefficient DICE, a metric used to evaluate the segmentation outcome. It is calculated using the formula $\text{DICE}=\frac{2|X\cap Y|}{|X|+|Y|}$, where $|X|$ and $|Y|$ denote the pixel counts in the segmentation and ground truth, respectively. Moreover, the criterion “DICE Bound” refers to the DICE coefficient of the segmentation achieving a value of $0.985$, indicating a high degree of similarity.
(a) Clover
(b) Label
(c) result 1
(d) result 2
(e) result 3
Figure 2: The performance of segmentation assignment.
In Figure 2, we present the segmentation results for an illustrative case: figure 2a is the original input image, figure 2b labels the prior that the white pixels label the positive prior, and the black pixels label the negative prior, while the remaining figures show the segmentation results under three different criteria (DICE Bound, $\|\nabla E(x)\|\leq 10^{-11}$ and $\|x^{n}-x^{n-1}\|\leq 10^{-11}$).
[…]
Algorithm
DICE Bound Iter 88 54 1003 341 191 53 58 39
Time(s) 45.67 29.34 76.72 38.12 19.81 7.72 8.56 7.83
I $10^{-1}$ Iter 119 80 1052 442 238 72 160 55
Time(s) 59.67 41.12 80.72 43.12 21.81 10.72 16.06 9.43
$10^{-3}$ Iter 315 174 3182 1284 739 231 291 115
Time(s) 137.48 91.36 237.24 126.04 65.40 23.86 26.33 13.54
$10^{-5}$ Iter 520 290 4891 2154 1201 399 411 197
Time(s) 218.65 151.98 92.3 192.1 60.6 41.18 65.5 21.19
$10^{-7}$ Iter 720 462 Max 3277 1801 550 554 272
Time(s) 298.64 241.24 291.58 153.98 60.65 50.06 27.84
$10^{-9}$ Iter 1001 600 Max 4641 2601 987 741 395
Time(s) 416.17 316.12 408.92 216.26 106.15 68.41 38.45
$10^{-11}$ Iter 1320 655 Max Max 3201 1203 931 589
Time(s) 537.25 345.42 254.27 126.37 86.24 54.79
II $10^{-1}$ Iter 136 78 1188 485 270 82 165 85
Time(s) 67.52 40.73 89.09 46.86 24.36 11.28 16.47 11.66
$10^{-3}$ Iter 325 167 3535 1393 771 253 285 114
Time(s) 141.35 87.76 237.24 126.04 65.4 26.86 26.33 13.89
$10^{-5}$ Iter 562 271 Max 2672 1399 409 405 198
Time(s) 235.65 141.43 239.17 118.97 42.78 36.33 22.21
$10^{-7}$ Iter 849 409 Max 3445 2117 755 497 278
Time(s) 350.93 213.45 306.12 181.71 79.58 46.47 28.31
$10^{-9}$ Iter 1099 556 Max 4656 2819 1084 687 356
Time(s) 449.59 291.34 410.21 230.09 115.96 62.01 35.39
$10^{-11}$ Iter 1379 670 Max Max 3525 1232 887 508
Time(s) 561.77 353.76 274.84 131.98 81.72 48.82
6 Conclusion
We investigate two kinds of proximal DC algorithms and their corresponding extrapolation, which are commonly used in non-convex optimization. The extrapolation parameter is completely determined by a non-monotone line search method. We prove the global convergence of the improved algorithm with the KL property. The numerical experiments show that the proposed algorithms are highly efficient. In addition, our non-monotone line search can also be performed using both methods from [12] and [23], which are also discussed in [15].
Acknowledgements Ran Zhang and Hongpeng Sun acknowledge the support of the National Natural Science Foundation of China under grant No.  12271521, National Key R&D Program of China (2022ZD0116800), and Beijing Natural Science Foundation No. Z210001. The second author acknowledges Prof. Ting Kei Pong of The Hong Kong Polytechnic University for suggesting the techniques developed in [22] during a private communication.